# On the arithmetic and geometry of spaces $L_{m+1,n}$

### v1

@@ 5.2. Subspaces of standard spaces via étale pullbacks
with non-zero leading coefficients. Let $Q_{11},\dots,Q_{n1}$ be a $n$-uple of polynomials giving rise to a basis of $\Omega_{1}$. Then by Proposition 4.17, up to multiplying by a constant, $Q_{11}(S_{1}(X)),\dots,Q_{n1}(S_{1}(X))$ give rise to a basis of $\Omega$. Similarly, if $Q_{12},\dots,Q_{n2}$ give rise to a basis of $\Omega_{2}$ then (up to multiplying by a constant) $Q_{12}(S_{2}(X)),\dots,Q_{n2}(S_{2}(X))$ give rise to a basis of $\Omega$, as well. Up to changing the basis of $\Omega_{2}$, we can assume that the bases of $\Omega$ produced in the two cases are the same, and then apply Proposition 4.1.ii to get that $Q_{j1}(S_{1}(X))=\eta Q_{j2}(S_{2}(X))$ for every $j=1,\dots,n$ and $\eta\in k^{\times}$. By definition we have $\deg(Q_{j1})=\lambda_{1}$ and $\deg(Q_{j2})=\lambda_{2}$, and the condition $\lambda_{1}<\lambda_{2}$ implies that $d_{1}>d_{2}$, since we ought to have $\lambda_{1}d_{1}=\lambda_{2}d_{2}$.
@@ 5.2.1. Standard spaces for $p=2$
Conversely, for every $n$-uple of elements $W_{1},\dots,W_{n}\in k^{\times}$ such that $\Delta_{n}(\underline{W})\neq 0$ one can set $V_{i}=\Delta_{n-1}(\underline{\hat{W_{i}}})\in k^{\times}$ and $U_{i}=\frac{V_{i}^{2}}{\Delta_{n}(\underline{W})}\in k^{\times}$, and show that the resulting
@@ 6. Classification of spaces $L_{12,2}$ and $L_{15,2}$ over ${\mathbb{F}}_{3}$
The aim of this section is to completely classify spaces $L_{12,2}$ and $L_{15,2}$ up to equivalence and Frobenius equivalence in the case where $p=3$. We recall that the spaces $L_{3,2},L_{6,2}$ and $L_{9,2}$ ($p=3$) are classified by Pagot in [8], so the results of this section are a natural prosecution of that work. By exhibiting the existence of a space $L_{15,2}$, we provide the first known example of a space $L_{\lambda p,2}$ where $p-1$ does not divide $\lambda$. This section relies on computations of Gröbner bases to solve polynomial systems in characteristic 3. The supporting Macaulay2 code can be found in a public repository555available at the url https://github.com/DanieleTurchetti/equidistant, in a form that can be easily replicated using other computer algebra systems.
@@ 6.1. Classification of spaces $L_{12,2}$
$Q_{1}:=a(X^{4}-s_{1}X^{3}+s_{2}X^{2}-s_{3}X+s_{4})$
$Q_{2}:=X^{4}-t_{1}X^{3}+t_{2}X^{2}-t_{3}X+t_{4},$
and we look for conditions such that the pair $(Q_{1},Q_{2})$ is a prompt for a space $L_{12,2}$. To this aim, we consider the expressions $R_{k}$’s of Convention 3.5 as polynomials with coefficients in $\{a,s_{1},\dots,s_{4},t_{1},\dots,t_{4}\}$ and look for a solution to the system of equations
@@ 2. $s_{3}=t_{3}=0$’:
      Suppose that this is not the case. Then, up to replacing $Q_{1}$ with $aQ_{2}$ and $Q_{2}$ with $\frac{1}{a}Q_{1}$, we can suppose that $s_{3}\neq 0$. Let $\alpha\in k$ such that $\alpha^{3}=s_{3}^{-1}$ and apply the transformation $X\to\alpha X$ to get that $s_{3}=1$. Then one finds that the system (6.1) has no solution (see [12, Program 6.1]), giving rise to a contradiction.
@@ 3. $s_{4}=a^{8}$, $t_{4}=1$’:
      A computation of Gröbner basis under the reductions above (see [12, Program 6.2]) returns the condition $s_{4}=a^{8}t_{4}$. Since $t_{3}=0$ we know that $t_{4}\neq 0$ otherwise the polynomial $B$ would have multiple roots. We can then pick $\beta\in k$ such that $\beta^{4}=t_{4}^{-1}$ and apply the transformation $X\to\beta X$ to get that $t_{4}=1$.
@@
    $Q_{1}:=a(X^{4}+s_{2}X^{2}+a^{8})$
    $Q_{2}:=X^{4}+t_{2}X^{2}+1.$
    Let us simplify the notation by setting $s:=s_{2}$ and $t:=t_{2}$, and compute these coefficients. Now, [12, Program 6.3] can compute a Grobner basis of the system (6.1). By inspecting the first element of the basis, we conclude that we have at most six possible expressions for $s$:
@@
    Let us show how the situation can be further simplified: first, we note that the transformation $X\to iX$ with $i^{2}=-1$ produces a new pair $(S_{1},S_{2})$ with
    $S_{1}:=a(X^{4}-s_{2}X^{2}+a^{8})$
    $S_{2}:=X^{4}-t_{2}X^{2}+1.$
    Then, we note that the pair $(-Q_{1},Q_{2})$ produces the same space $L_{12,2}$. Hence, if we can find a solution to the system (6.1) for a pair $(a,s)$ we also have a solution for pairs $(-a,s)$,$(a,-s)$, and $(-a,-s)$, which reduces our search to the following two situations:
@@ 1. Case 1:
      $s=a^{4}-a^{2}-1$. In this case, Program [12, Program 6.4] returns $t=-(a^{4}+a^{2}-1)$ (under the condition that $s\neq 0$), and then we find that $Q_{1}=Q_{1,a}$ and $Q_{2}=Q_{2,a}$, as required. If $s=0$, then the program returns $t^{2}=a^{2}+1=a^{4}$, which gives the two values $t=\pm a^{2}=\pm(a^{4}+a^{2}-1)$. Note that these two values correspond to equivalent spaces, under the transformation $X\mapsto iX$. We conclude that, in Case 1, the pair $(Q_{1},Q_{2})$ is a prompt for a space equivalent to $\Omega_{a}$.
@@ 2. Case 2:
      $s=a^{4}+a^{2}+a+1$. In this case, Program [12, Program 6.5] returns $t=-(a^{4}-a^{3}+a^{2}+1)$ (under the condition that $s\neq 0$).
@@
$f_{1}=\prod_{\begin{subarray}{c}\epsilon_{1}=\pm 1\\ \epsilon_{2}=\pm 1\end{subarray}}\left(X-(\epsilon_{1}(a-1)j+\epsilon_{2}ai)\right)^{\epsilon_{1}}\prod_{\begin{subarray}{c}\epsilon_{1}=\pm 1\\ \epsilon_{2}=\pm 1\end{subarray}}\left(X-(\epsilon_{1}(a+1)j+\epsilon_{2}ai)\right)^{\epsilon_{1}}\prod_{\begin{subarray}{c}\epsilon_{1}=\pm 1\\ \epsilon_{2}=\pm 1\end{subarray}}\left(X-(\epsilon_{1}aj+\epsilon_{2}(a^{2}-1))\right)^{\epsilon_{1}}$
@@
$f_{2}=\prod_{\begin{subarray}{c}\epsilon_{1}=\pm 1\\ \epsilon_{2}=\pm 1\end{subarray}}\left(X-(\epsilon_{1}(a-1)j+\epsilon_{2}ai)\right)^{-\epsilon_{1}}\prod_{\begin{subarray}{c}\epsilon_{1}=\pm 1\\ \epsilon_{2}=\pm 1\end{subarray}}\left(X-(\epsilon_{1}(a+1)j+\epsilon_{2}ai)\right)^{\epsilon_{1}}\prod_{\begin{subarray}{c}\epsilon_{1}=\pm 1\\ \epsilon_{2}=\pm 1\end{subarray}}\left(X-(\epsilon_{1}j+\epsilon_{2}(a^{2}-1))\right)^{\epsilon_{1}}.$
@@
$Z(Q_{2,a})=\{aj+(a^{2}-1),aj-(a^{2}-1),-aj+(a^{2}-1),-aj-(a^{2}-1)\},$
as expected. By Theorem 6.2 and Remark 3.4 we know that $\Omega_{a}$ is generated by the differential forms $\omega_{1}:=\frac{Q_{1,a}dX}{c^{3}(Q_{1,a}^{3}Q_{2,a}-Q_{1,a}Q_{2,a}^{3})}$ and $\omega_{2}:=\frac{Q_{2,a}dX}{c^{3}(Q_{1,a}^{3}Q_{2,a}-Q_{1,a}Q_{2,a}^{3})}$ with $c^{6}(a^{2}+1)^{5}(a^{3}-a)^{10}=1$. To compute the residues of these forms at their poles, we remark that $c^{3}=\pm\frac{1}{(a^{2}+1)^{5/2}(a^{3}-a)^{5}}=\pm\frac{1}{j^{5}(a^{3}-a)^{5}}$. For each $x\in Z(Q_{1,a})\cup Z(Q_{1,a}+Q_{2,a})\cup Z(Q_{1,a}-Q_{2,a})\cup Z(Q_{2,a})$ we can compute the quantities $(Q_{1,a}^{2}Q_{2,a}-Q_{2,a}^{3})^{\prime}(x)$ and $(Q_{1,a}^{3}-Q_{1,a}Q_{2,a}^{2})^{\prime}(x)$ and check when they are equal to ±$\pm j^{5}(a^{3}-a)^{5}$. Up to possibly replacing $j$ with $-j$, and using the formulas $res_{1}(x)=\frac{1}{c^{3}(Q_{1,a}^{2}Q_{2,a}-Q_{2,a}^{3})^{\prime}(x)}$ and $res_{2}(x)=\frac{1}{c^{3}(Q_{1,a}^{3}-Q_{1,a}Q_{2,a}^{2})^{\prime}(x)}$, we get the table of residues above, or equivalently that $\omega_{1}=\frac{df_{1}}{f_{1}}$ and $\omega_{2}=\frac{df_{2}}{f_{2}}$ with $f_{1},f_{2}$ as in the statement. ∎
@@ 6.2. Classification of spaces $L_{15,2}$
$Q_{1}(X):=(\mu^{2}-\mu-1)P_{0}(X)\;\;\text{and}\;\;Q_{2}(X):=-\mu P_{\infty}(X).$
One verifies that the second derivative of the polynomial $(Q_{1}^{3}-Q_{1}Q_{2}^{2})^{2}$ is equal to $-1$. Moreover, since $(\mu^{2}-\mu-1)\pm(-\mu)\notin{\mathbb{F}}_{3}$, the degree of $iQ_{1}+jQ_{2}$ is equal to 4 for every $[i:j]\in{\mathbb{P}}^{1}({\mathbb{F}}_{3})$. Then, we can set $\omega_{1}=\frac{Q_{1}}{Q_{1}^{3}Q_{2}-Q_{1}Q_{2}^{3}}dX$ and $\omega_{2}=\frac{Q_{2}}{Q_{1}^{3}Q_{2}-Q_{1}Q_{2}^{3}}dX$ and apply Proposition 3.1 to show that $\Omega_{1}:=\langle\omega_{1},\omega_{2}\rangle$ is a ${\mathbb{F}}_{3}$-vector space $L_{15,2}$.
To write the generators of this spaces in logarithmic form, we need to compute their residues at the poles. In order to do this, we denote by $Z_{1}$ the set of zeroes of $Q_{1}^{2}Q_{2}-Q_{2}^{3}$, by $Z_{2}$ the set of zeroes of $Q_{1}^{3}-Q_{1}Q_{2}^{2}$, and we introduce the functions $res_{1}:Z_{1}\to{\mathbb{F}}_{3}^{\times}$ and $res_{2}(x):Z_{2}\to{\mathbb{F}}_{3}^{\times}$ defined by $res_{i}(x):=res_{x}(\omega_{i})$. Then, we compute them thanks to the formulas $res_{1}(x)=\frac{1}{(Q_{1}^{2}Q_{2}-Q_{2}^{3})^{\prime}(x)}$ and $res_{2}(x)=\frac{1}{(Q_{1}^{3}-Q_{1}Q_{2}^{2})^{\prime}(x)}$ We find the following table of values:
$x$ $res_{1}(x)$ $res_{2}(x)$
$\mu^{4}$ 0 1
$\mu^{9}$ 0 1
$\mu^{19}$ 0 1
$\mu$ 0 $-1$
00 0 $-1$
$\mu^{5}$ $-1$ 1
$\mu^{15}$ 1 $-1$
$\mu^{24}$ 1 $-1$
$\mu^{8}$ 1 $-1$
$\mu^{21}$ 1 $-1$
$x$ $res_{1}(x)$ $res_{2}(x)$
$\mu^{25}$ $-1$ $-1$
$\mu^{13}$ $1$ $1$
$\mu^{18}$ $-1$ $-1$
$\mu^{12}$ $1$ $1$
$\mu^{17}$ $-1$ $-1$
$\mu^{10}$ $1$ 0
$\mu^{3}$ $1$ 0
$\mu^{2}$ $-1$ 0
$\mu^{7}$ $-1$ 0
$\mu^{22}$ $1$ 0
@@
[…]
@@
$Q_{1}(X):=aP_{0}(X)\;\;\text{and}\;\;Q_{2}(X):=-(a+1)P_{\infty}(X),$
and we can verify that the second derivative of the polynomial $(Q_{1}^{3}-Q_{1}Q_{2}^{2})^{2}$ is equal to $-1$. Moreover, the degree of $iQ_{1}+jQ_{2}$ is equal to 5 for every $[i:j]\in{\mathbb{P}}^{1}({\mathbb{F}}_{3})$. Hence, if we set $\omega_{1}=\frac{Q_{1}}{Q_{1}^{3}Q_{2}-Q_{1}Q_{2}^{3}}dX$ and $\omega_{2}=\frac{Q_{2}}{Q_{1}^{3}Q_{2}-Q_{1}Q_{2}^{3}}dX$ and apply Proposition 3.1 this shows that $\Omega_{2}:=\langle\omega_{1},\omega_{2}\rangle$ is a space $L_{15,2}$.
As in Example 6.5, we can easily compute the residues of $\omega_{1}$ and $\omega_{2}$ at their poles, and get the following table.
$x$ $res_{1}(x)$ $res_{2}(x)$
$\mu^{7}$ 0 $-1$
$\mu^{30}$ 0 $-1$
$\mu^{51}$ 0 $-1$
$\mu^{59}$ 0 $-1$
$\mu^{63}$ 0 $-1$
$\mu^{26}$ $1$ $-1$
$\mu^{50}$ $-1$ $1$
$\mu^{52}$ $-1$ $1$
$\mu^{68}$ $-1$ $1$
$\mu^{74}$ 1 $-1$
$x$ $res_{1}(x)$ $res_{2}(x)$
$\mu^{34}$ $-1$ $-1$
$\mu^{60}$ $1$ $1$
$\mu^{66}$ $-1$ $-1$
$\mu^{70}$ $1$ $1$
00 $1$ $1$
$\mu^{10}$ $1$ 0
$\mu^{11}$ $1$ 0
$\mu^{19}$ $1$ 0
$\mu^{20}$ $1$ 0
$\mu^{40}$ $-1$ 0
As in the previous example we have $\omega_{1}=\frac{dF_{1}}{F_{1}}$ and $\omega_{2}=\frac{dF_{2}}{F_{2}}$ with
@@ 00 $1$ $1$
Our main result of this section says that the examples above are essentially the only spaces $L_{15,2}$.
@@
$Q_{2}:=X^{5}-t_{1}X^{4}+t_{2}X^{3}-t_{3}X^{2}+t_{4}X-t_{5}.$
Then, we consider the expressions $R_{k}$’s of Convention 3.5 as polynomials with coefficients in $\{a,s_{1},\dots,s_{5},t_{1},\dots,t_{5}\}$ and we aim to find a solution to the system of equations
@@
The existence of a space $L_{15,2}$ gives rise to a solution to the system (6.8) (see Convention 3.5). Conversely a solution to (6.8) produces a space $L_{15,2}$. In fact, such a solution would imply that the second derivative of $(Q_{1}^{3}-Q_{1}Q_{2}^{2})^{2}$ is equal to the nonzero constant $-R_{1}(a,s_{i},t_{i})$. Then, Proposition 3.1 and Remark 3.4 tell us that we can build a space $L_{15,2}$ from $Q_{1}$ and $Q_{2}$ by setting $\omega_{1}=\frac{Q_{1}}{Q_{1}^{3}Q_{2}-Q_{1}Q_{2}^{3}}dX$ and $\omega_{2}=\frac{Q_{2}}{Q_{1}^{3}Q_{2}-Q_{1}Q_{2}^{3}}dX$, and considering the space $\langle\omega_{1},\omega_{2}\rangle$.
Unfortunately, solving (6.8) is not an easy task even when assisted by a computer (a brute-force calculation of Gröbner basis turns out to be hopeless without further assumptions), so we need to simplify the equations before we go further with our strategy. To this aim, we apply Lemma 3.6, we set $s_{1}=t_{1}$ and we reparametrize ${\mathbb{P}}^{1}_{k}$ in such a way to have $s_{1}=t_{1}=0$. Then, we obtain the following result.

### v2

@@ 5.2. Subspaces of standard spaces via étale pullbacks
with non-zero leading coefficients. Let $Q_{11},\dots,Q_{n1}$ be a $n$-tuple of polynomials giving rise to a basis of $\Omega_{1}$. Then by Proposition 4.19, up to multiplying by a constant, $Q_{11}(S_{1}(X)),\dots,Q_{n1}(S_{1}(X))$ give rise to a basis of $\Omega$. Similarly, if $Q_{12},\dots,Q_{n2}$ give rise to a basis of $\Omega_{2}$ then (up to multiplying by a constant) $Q_{12}(S_{2}(X)),\dots,Q_{n2}(S_{2}(X))$ give rise to a basis of $\Omega$, as well. Up to changing the basis of $\Omega_{2}$, we can assume that the bases of $\Omega$ produced in the two cases are the same, and then apply Proposition 4.1.ii to get that $Q_{j1}(S_{1}(X))=\eta Q_{j2}(S_{2}(X))$ for every $j=1,\dots,n$ and $\eta\in k^{\times}$. By definition we have $\deg(Q_{j1})=\lambda_{1}$ and $\deg(Q_{j2})=\lambda_{2}$, and the condition $\lambda_{1}<\lambda_{2}$ implies that $d_{1}>d_{2}$, since we ought to have $\lambda_{1}d_{1}=\lambda_{2}d_{2}$.
@@ 5.2.1. Standard spaces for $p=2$
Conversely, for every $n$-tuple of elements $W_{1},\dots,W_{n}\in k^{\times}$ such that $\Delta_{n}(\underline{W})\neq 0$ one can set $V_{i}=\Delta_{n-1}(\underline{\hat{W_{i}}})\in k^{\times}$ and $U_{i}=\frac{V_{i}^{2}}{\Delta_{n}(\underline{W})}\in k^{\times}$, and show that the resulting
@@ 6. Classification of spaces $L_{12,2}$ and $L_{15,2}$ over ${\mathbb{F}}_{3}$
The aim of this section is to completely classify spaces $L_{12,2}$ and $L_{15,2}$ up to equivalence and Frobenius equivalence in the case where $p=3$. We recall that the spaces $L_{3,2},L_{6,2}$ and $L_{9,2}$ ($p=3$) are classified by Pagot in [Pagot02], so the results of this section are a natural prosecution of that work. By exhibiting the existence of a space $L_{15,2}$, we provide the first known example of a space $L_{\lambda p,2}$ where p1p-1 does not divide $\lambda$. This section relies on computations of Gröbner bases to solve polynomial systems in characteristic 3. The supporting Macaulay2 code can be found in a public repository444available at the url https://github.com/DanieleTurchetti/equidistant, in a form that can be easily replicated using other computer algebra systems.
@@ 6.1. Classification of spaces $L_{12,2}$
$Q_{1}:=X^{4}-t_{1}X^{3}+t_{2}X^{2}-t_{3}X+t_{4},$
$Q_{2}:=a(X^{4}-s_{1}X^{3}+s_{2}X^{2}-s_{3}X+s_{4}),$
and we look for conditions such that the pair $(Q_{1},Q_{2})$ is a prompt for a space $L_{12,2}$. To this aim, we consider the expressions $R_{k}$’s of Convention 3.7 as polynomials with coefficients in $\{a,s_{1},\dots,s_{4},t_{1},\dots,t_{4}\}$ and look for a solution to the system of equations
@@ 2. $s_{3}=t_{3}=0$’:
      Suppose by contradiction that this is not the case. Then, up to replacing $Q_{1}$ with $\frac{Q_{2}}{a}$ and $Q_{2}$ with $aQ_{1}$, we can assume that $s_{3}\neq 0$. In fact, after applying a homothety, we can assume that $s_{3}=1$. Then by resolving system (6.1) computationally, one finds that it has no solution (see [GithubRepo, Program 6.1]).
@@ 3. $s_{4}=a^{8}$, $t_{4}=1$’:
      A computation of Gröbner basis under the reductions above (see [GithubRepo, Program 6.2]) returns the condition $s_{4}=a^{8}t_{4}$. Since $t_{3}=0$ we know that $t_{4}\neq 0$ otherwise the polynomial $B$ would have multiple roots. We can then pick $\beta\in k$ such that $\beta^{4}=t_{4}^{-1}$ and apply the transformation $X\to\beta X$ to get that $t_{4}=1$.
@@
    $Q_{1}:=X^{4}+t_{2}X^{2}+1,$
    $Q_{2}:=a(X^{4}+s_{2}X^{2}+a^{8}).$
    Let us simplify the notation by setting $s:=s_{2}$ and $t:=t_{2}$, and compute these coefficients, again using Gröbner bases (see [GithubRepo, Program 6.3]). By inspecting the first element of the basis, we conclude that we have at most six possible expressions for $s$:
@@
    Let us show how the situation can be further simplified: first, we note that the transformation $X\to iX$ with $i^{2}=-1$ sends $(Q_{1},Q_{2})$ to the pair of polynomials
    $\left(X^{4}-tX^{2}+1,a(X^{4}-sX^{2}+a^{8})\right).$
    Further, we note that $(Q_{1},Q_{2})$ is a prompt for a space $L_{12,2}$ if, and only if, the pair $(-Q_{1},Q_{2})$ is a prompt for the same space. As a result of these two observations, if we can find a solution to the system (6.1) for a pair $(a,s)$ we also have a solution for pairs (a,s)(-a,s),(a,s)(a,-s), and (a,s)(-a,-s), which reduces our search to the following two situations:
@@ 1. Case 1:
      $s=a^{4}-a^{2}-1$. In this case, Program [GithubRepo, Program 6.4] returns $t=-(a^{4}+a^{2}-1)$ (under the condition that $s\neq 0$), and then we find that $Q_{1}=Q_{1,a}$ and $Q_{2}=Q_{2,a}$, as required. If $s=0$, then the program returns $t^{2}=a^{2}+1=a^{4}$, which gives the two values $t=\pm a^{2}=\pm(a^{4}+a^{2}-1)$. Note that these two values correspond to equivalent spaces, under the transformation $X\mapsto iX$. We conclude that, in Case 1, the pair $(Q_{1},Q_{2})$ is a prompt for a space equivalent to $\Omega_{a}$.
@@ 2. Case 2:
      $s=a^{4}+a^{2}+a+1$. In this case, Program [GithubRepo, Program 6.5] returns $t=-(a^{4}-a^{3}+a^{2}+1)$ (under the condition that $s\neq 0$).
@@
$f_{1}=\prod_{\begin{subarray}{c}\epsilon_{1}=\pm 1\\ \epsilon_{2}=\pm 1\end{subarray}}\left(X-(\epsilon_{1}(a-1)j+\epsilon_{2}ai)\right)^{-\epsilon_{1}}\prod_{\begin{subarray}{c}\epsilon_{1}=\pm 1\\ \epsilon_{2}=\pm 1\end{subarray}}\left(X-(\epsilon_{1}(a+1)j+\epsilon_{2}ai)\right)^{\epsilon_{1}}\prod_{\begin{subarray}{c}\epsilon_{1}=\pm 1\\ \epsilon_{2}=\pm 1\end{subarray}}\left(X-(\epsilon_{1}j+\epsilon_{2}(a^{2}-1))\right)^{\epsilon_{1}}$
@@
$f_{2}=\prod_{\begin{subarray}{c}\epsilon_{1}=\pm 1\\ \epsilon_{2}=\pm 1\end{subarray}}\left(X-(\epsilon_{1}(a-1)j+\epsilon_{2}ai)\right)^{\epsilon_{1}}\prod_{\begin{subarray}{c}\epsilon_{1}=\pm 1\\ \epsilon_{2}=\pm 1\end{subarray}}\left(X-(\epsilon_{1}(a+1)j+\epsilon_{2}ai)\right)^{\epsilon_{1}}\prod_{\begin{subarray}{c}\epsilon_{1}=\pm 1\\ \epsilon_{2}=\pm 1\end{subarray}}\left(X-(\epsilon_{1}aj+\epsilon_{2}(a^{2}-1))\right)^{\epsilon_{1}}.$
@@
$Z(Q_{2,a})=\{j+a^{2}-1,j-a^{2}+1,-j+a^{2}-1,-j-a^{2}+1\},$
as expected. By Theorem 6.2 and Remark 3.6 we know that $\Omega_{a}$ is generated by the differential forms $\omega_{1}:=\frac{Q_{2,a}dX}{c^{3}(Q_{1,a}^{3}Q_{2,a}-Q_{1,a}Q_{2,a}^{3})}$ and $\omega_{2}:=\frac{-Q_{1,a}dX}{c^{3}(Q_{1,a}^{3}Q_{2,a}-Q_{1,a}Q_{2,a}^{3})}$ with $c^{6}(a^{2}+1)^{5}(a^{3}-a)^{10}=1$. To compute the residues of these forms at their poles, we remark that $c^{3}=\pm\frac{1}{(a^{2}+1)^{5/2}(a^{3}-a)^{5}}=\pm\frac{1}{j^{5}(a^{3}-a)^{5}}$. For each $x\in Z(Q_{1,a})\cup Z(Q_{1,a}+Q_{2,a})\cup Z(Q_{1,a}-Q_{2,a})\cup Z(Q_{2,a})$ we can compute the quantities $(Q_{1,a}^{2}Q_{2,a}-Q_{2,a}^{3})^{\prime}(x)$ and $(Q_{1,a}^{3}-Q_{1,a}Q_{2,a}^{2})^{\prime}(x)$ and check when they are equal to $\pm j^{5}(a^{3}-a)^{5}$. Up to possibly replacing $j$ with j-j, and using the formulas $\mathrm{res}_{1}(x)=\frac{1}{c^{3}(Q_{1,a}^{3}-Q_{1,a}Q_{2,a}^{2})^{\prime}(x)}$ and $\mathrm{res}_{2}(x)=\frac{1}{c^{3}(Q_{1,a}^{2}Q_{2,a}-Q_{2,a}^{3})^{\prime}(x)}$ we get the table of residues above, or equivalently that $\omega_{1}=\frac{df_{1}}{f_{1}}$ and $\omega_{2}=\frac{df_{2}}{f_{2}}$ with $f_{1},f_{2}$ as in the statement. ∎
@@ 6.2. Classification of spaces $L_{15,2}$
$Q_{1}(X):=-\mu P_{\infty}(X)\;\;\text{and}\;\;Q_{2}(X):=(\mu^{2}-\mu-1)P_{0}(X).$
One verifies that the second derivative of the polynomial $(Q_{1}^{3}-Q_{1}Q_{2}^{2})^{2}$ is equal to 1-1. Moreover, since $(\mu^{2}-\mu-1)\pm(-\mu)\notin{\mathbb{F}}_{3}$, the degree of $iQ_{1}+jQ_{2}$ is equal to 5 for every $i:j]\in{\mathbb{P}}^{1}({\mathbb{F}}_{3})$. Then, we can set $\omega_{1}=\frac{Q_{2}}{Q_{1}^{3}Q_{2}-Q_{1}Q_{2}^{3}}dX$ and $\omega_{2}=\frac{-Q_{1}}{Q_{1}^{3}Q_{2}-Q_{1}Q_{2}^{3}}dX$ and apply Proposition 3.1 to show that $\Omega_{1}:=\langle\omega_{1},\omega_{2}\rangle$ is a ${\mathbb{F}}_{3}$-vector space $L_{15,2}$.
To write the generators of this spaces in logarithmic form, we need to compute their residues at the poles. In order to do this, we denote by $Z_{1}$ the set of zeroes of $Q_{1}^{3}-Q_{1}Q_{2}^{2}$, by $Z_{2}$ the set of zeroes of $Q_{1}^{2}Q_{2}-Q_{2}^{3}$, and we introduce the functions $\mathrm{res}_{1}:Z_{1}\to{\mathbb{F}}_{3}^{\times}$ and $\mathrm{res}_{2}(x):Z_{2}\to{\mathbb{F}}_{3}^{\times}$ defined by $\mathrm{res}_{i}(x):=\mathrm{res}_{x}(\omega_{i})$. Then, we compute them thanks to the formulas $\mathrm{res}_{1}(x)=\frac{1}{(Q_{1}^{3}-Q_{1}Q_{2}^{2})^{\prime}(x)}$ and $\mathrm{res}_{2}(x)=\frac{1}{(Q_{1}^{2}Q_{2}-Q_{2}^{3})^{\prime}(x)}$. We find the following table of values:
$x$ $\mathrm{res}_{1}(x)$ $\mathrm{res}_{2}(x)$
$\mu^{4}$ 1 0
$\mu^{9}$ 1 0
$\mu^{19}$ 1 0
$\mu$ 1-1 0
0 1-1 0
$\mu^{5}$ 11 -1
$\mu^{15}$ 1-1 11
$\mu^{24}$ 1-1 11
$\mu^{8}$ 1-1 11
$\mu^{21}$ 1-1 11
$x$ $\mathrm{res}_{1}(x)$ $\mathrm{res}_{2}(x)$
$\mu^{25}$ 1-1 1-1
$\mu^{13}$ 11 11
$\mu^{18}$ 1-1 1-1
$\mu^{12}$ 11 11
$\mu^{17}$ 1-1 1-1
$\mu^{10}$ 0 11
$\mu^{3}$ 0 11
$\mu^{2}$ 0 1-1
$\mu^{7}$ 0 1-1
$\mu^{22}$ 0 1
@@
Let ${\mathbb{F}}_{81}$ be the finite field with 81 elements, and let us write ${\mathbb{F}}_{81}={\mathbb{F}}_{3}[\mu]$, with $\mu^{4}+\mu^{3}-\mu^{2}-\mu-1=0$. [^5] This is equivalent to write ${\mathbb{F}}_{81}={\mathbb{F}}_{3}[\alpha]$ for a choice of $\alpha$ satisfying $\alpha^{4}+\alpha^{2}-1=0$ and set $\mu=\alpha^{3}+\alpha-1$. We have that $\mu^{40}=-1$, and then $\mu$ is a generator of the cyclic multiplicative group ${\mathbb{F}}_{81}^{\times}$. In analogy with Example 6.5, we define for every element of ${\mathbb{P}}^{1}({\mathbb{F}}_{3})$ a subset of ${\mathbb{F}}_{81}$ as follows:
@@
$Q_{1}(X):=-(a+1)P_{\infty}(X)\;\;\text{and}\;\;Q_{2}(X):=aP_{0}(X),$
and we can verify that the second derivative of the polynomial $(Q_{1}^{3}-Q_{1}Q_{2}^{2})^{2}$ is equal to 1-1. Moreover, the degree of $iQ_{1}+jQ_{2}$ is equal to 5 for every $i:j]\in{\mathbb{P}}^{1}({\mathbb{F}}_{3})$. Hence, if we set $\omega_{1}=\frac{Q_{2}}{Q_{1}^{3}Q_{2}-Q_{1}Q_{2}^{3}}dX$ and $\omega_{2}=\frac{-Q_{1}}{Q_{1}^{3}Q_{2}-Q_{1}Q_{2}^{3}}dX$ and apply Proposition 3.1 this shows that $\Omega_{2}:=\langle\omega_{1},\omega_{2}\rangle$ is a space $L_{15,2}$.
As in Example 6.5, we can easily compute the residues of $\omega_{1}$ and $\omega_{2}$ at their poles, and get the following table
$x$ $\mathrm{res}_{1}(x)$ $\mathrm{res}_{2}(x)$
$\mu^{7}$ 1-1 0
$\mu^{30}$ 1-1 0
$\mu^{51}$ 1-1 0
$\mu^{59}$ 1-1 0
$\mu^{63}$ 1-1 0
$\mu^{26}$ 1-1 11
$\mu^{50}$ 11 1-1
$\mu^{52}$ 11 1-1
$\mu^{68}$ 11 1-1
$\mu^{74}$ 1-1 11
$x$ $\mathrm{res}_{1}(x)$ $\mathrm{res}_{2}(x)$
$\mu^{34}$ 1-1 1-1
$\mu^{60}$ 11 11
$\mu^{66}$ 1-1 1-1
$\mu^{70}$ 11 11
0 11 11
$\mu^{10}$ 0 1
$\mu^{11}$ 0 1
$\mu^{19}$ 0 1
$\mu^{20}$ 0 1
$\mu^{40}$ 0 1-1
and as in the previous example we have $\omega_{1}=\frac{dF_{1}}{F_{1}}$ and $\omega_{2}=\frac{dF_{2}}{F_{2}}$ with
@@
The main result of this section states that all the spaces $L_{15,2}$ can be obtained from the examples above.
@@
$Q_{2}:=a(X^{5}-s_{1}X^{4}+s_{2}X^{3}-s_{3}X^{2}+s_{4}X-s_{5}).$
Then, we consider the expressions $R_{k}$’s of Convention 3.7 as polynomials with coefficients in $\{a,s_{1},\dots,s_{5},t_{1},\dots,t_{5}\}$ and we aim to find a solution to the system of equations
@@
The existence of a space $L_{15,2}$ gives rise to a solution to the system (6.8) (see Convention 3.7). Conversely a solution to (6.8) produces a space $L_{15,2}$. In fact, such a solution would imply that the second derivative of $(Q_{1}^{3}-Q_{1}Q_{2}^{2})^{2}$ is equal to the nonzero constant $R_{1}(a,s_{i},t_{i})$. It follows that $(Q_{1},Q_{2})$ is a prompt for a space $L_{15,2}$.
Unfortunately, solving (6.8) is not an easy task even when assisted by a computer (a brute-force calculation of Gröbner basis turns out to be hopeless without further assumptions), so we need to simplify the equations before we go further with our strategy. To this aim, we apply Lemma 3.8 to get that $s_{1}=t_{1}$ and we reparametrize ${\mathbb{P}}^{1}_{k}$ in such a way to have $s_{1}=t_{1}=0$. Then, we obtain the following result.